OpenStudy (wolfmon):

(08.03)A set of equations is given below: Equation A: y = x + 1 Equation B: y = 4x + 5 Which of the following steps can be used to find the solution to the set of equations? x + 1 = 4x + 5 x = 4x + 5 x + 1 = 4x x + 5 = 4x + 1

OpenStudy (wolfmon):

Please i need help!!

OpenStudy (butterflydreamer):

looking at equation A and equation B, can you see that BOTH have values of y? " Equation A: y = x + 1 Equation B: y = 4x + 5 " See how both have "y= ..." This means that both "x + 1" and "4x + 5" are values of y :)

OpenStudy (wolfmon):

ohh ok thank you

OpenStudy (anonymous):

Hint: two things are equal to the same thing, are equal to each other!

OpenStudy (wolfmon):

i get it now thank you so much!

OpenStudy (butterflydreamer):

:) yaay you're welcome